OpenStudy (anonymous):

need help...first time on site: limit to inifinity of this n^.5 plus n^(1/3) divided by n + 2n^(2/3) book pulled out n to the neg .5 not sure why. Am thinking there is a better way? Don't you always divide by highest power?

OpenStudy (zehanz):

IMO, if you divide everything by n^(2/3), you get:\[\frac{ n^{-\frac{1}{6}} +n^{-\frac{1}{3}}}{ n^{\frac{1}{3}}+2 }\]Now, if n goes to infinity, the powers of n in the numerator go to 0, the denominator goes to infinity, so the whole thing goes to 0. So, I think you're right!

OpenStudy (anonymous):

 I believe you divide by the lowest power, which is 1/3. so it would be.